# On uniqueness of KP soliton structures

### v1

@@ 3. $\left(iii\right)$
The proof of (i) is a consequence of the following identity. One has
@@ 3.1. Simple Phases
As in the proof of Lemma 3.7, we only prove the necessary condition. Thanks to this last result, $\Theta\left(t,x,y\right)=A\left(t,x\right)\exp\left(c\left(t,x\right)y\right)$. Assuming now $W_{x}\left(\Theta\right)=0$, i.e. $\Theta\Theta_{xxxx}-\Theta^{2}_{xx}=0$, and replacing $\Theta$ in this equation, one gets the next degree-3 polynomial in the $y$ variable:
@@ 4.1. Airy and Heat structures
In the case $c_{2}\left(t\right)=0$ we first have $\Theta\left(t,x,y\right)=A\left(t,x\right)\exp\left(c_{1}\left(t\right)y\right)$. Since $H(\Theta)=0$ necessarily $A(t,x)c_{1}(t)-A_{xx}(t,x)=0$. Depending on the sign of $c_{1}(t)$, we have
$A(t,x)=a_{0}(t)e^{\sqrt{c_{1}(t)}x}+a_{1}(t)e^{-\sqrt{c_{1}(t)}x},\quad c_{1}(t)>0,$
$A(t,x)=a_{0}(t)+a_{1}(t)x,\quad c_{1}(t)=0,$
@@
$A(t,x)=a_{0}(t)\cos(\sqrt{-c_{1}(t)}x)+a_{1}(t)\sin(\sqrt{-c_{1}(t)}x),\quad-c_{1}(t)>0.$
Since $\Theta>0$ we discard the third case, and in the second case we get $c_{2}(t)=0$, leading to a solution included in the first case. Note that naturally the condition $A$ is of $W_{x}$ type is satisfied. Denoting $k_{1}:=\sqrt{c_{1}}$ we conclude. This proves (i).
@@
$\Theta=A\left(t,x\right)\exp\left(\left(c_{1}(t)+c_{2}\left(t\right)x\right)y\right).$
@@
Case $c_{2}\left(t\right)=0$. First, we will examine the case in which $c_{2}\left(t\right)=0$. In this case from Corollary 3.9 $\Theta=A\left(t,x\right)\exp\left(c_{1}\left(t\right)y\right)$, with $W_{x}(A)=0$. First we compute
@@
$\Theta\left(t,x,y\right)=a_{1}(t)\exp\left(k_{1}x+k_{1}^{2}y\right)+a_{2}(t)\exp\left(-k_{1}x+k_{1}^{2}y\right).$
@@
$(a_{1}^{\prime}-a_{1}k_{1}^{3})\exp\left(k_{1}x+k_{1}^{2}y\right)+(a_{2}^{\prime}+a_{2}k_{1}^{3})\exp\left(-k_{1}x+k_{1}^{2}y\right)=0.$
We obtain $a_{1}(t)=a_{1}e^{k_{1}^{3}t}$, $a_{2}(t)=a_{2}e^{-k_{1}^{3}t}$, where $a_{1}$ and $a_{2}$ are constants. Therefore,
@@
$\Theta\left(t,x,y\right)=\exp\left(y\partial_{x}^{2}\right)\Theta_{0}\left(t,x\right)=\frac{1}{\left(4\pi y\right)^{1/2}}\int_{\mathbb{R}}\exp\left(-\frac{\left(x-s\right)^{2}}{4\pi y}\right)\Theta_{0}\left(t,s\right)\mathrm{d}s$
@@ 4.2. Soliton structures
where $a_{1},a_{2}>0$, and $\theta_{j}:=k_{j}x+k_{j}^{2}y+k_{j}^{3}t$, $k_{1},k_{2}\in\mathbb{R}$. We shall say that $\Theta$ represents a line-soliton if $\Theta$ has the previous form. Notice that unless $F=\log$, one does not have that $u=2\partial_{x}^{2}F(\Theta)$ is the standard KP line-soliton (1.13).
@@
Assume that $\Theta\in\mathcal{W}_{M}$, that is for $0\leq k_{1}\left(t,x\right)<k_{2}\left(t,x\right)<\ldots<k_{N}\left(t,x\right)$,
@@
Note now that $\left(\sum^{M}_{n=1}a_{n}\sum^{M}_{n=1}b_{n}=\sum^{M}_{n=1}c_{n}\hbox{, where }c_{n}=\sum^{n}_{i=0}a_{i}b_{n-i}\right)$,
@@
$=\sum^{M}_{n=1}\sum^{n}_{i=1}k^{4}_{n-i+1}a_{i}a_{n-i+1}\exp\left(\theta_{i}+\theta_{n-i+1}\right)$
$\quad-\sum^{M}_{n=1}\sum^{n}_{i=1}k^{2}_{i}k^{2}_{n-i+1}a_{i}a_{n-i+1}\exp\left(\theta_{i}+\theta_{n-i}\right)$
$=\sum^{M}_{n=1}\sum^{n}_{i=1}a_{i}a_{n-i+1}\left(k^{4}_{n-i+1}-k^{2}_{i}k^{2}_{n-i+1}\right)\exp\left(\theta_{i}+\theta_{n-i+1}\right).$
Since $\theta_{i}+\theta_{n-i+1}=\theta_{n-i+1}+\theta_{i}$,
$\Theta W_{y}\left(\Theta\right)$ $=\sum^{M}_{n=1}\sum^{n}_{i=1}a_{i}a_{n-i+1}\left(k^{4}_{n-i+1}-k^{2}_{i}k^{2}_{n-i+1}\right)\exp\left(\theta_{i}+\theta_{n-i+1}\right)$
$=\frac{1}{2}\sum^{M}_{n=1}\sum^{n}_{i=1}a_{i}a_{n-i+1}\left(k^{4}_{n-i+1}-k^{2}_{i}k^{2}_{n-i+1}+k^{4}_{i}-k^{2}_{n-i+1}k^{2}_{i}\right)\exp\left(\theta_{i}+\theta_{n-i+1}\right)$
$=\frac{1}{2}\sum^{M}_{n=1}\sum^{n}_{i=1}a_{i}a_{n-i+1}\left(k^{4}_{n-i+1}-2k^{2}_{i}k^{2}_{n-i+1}+k^{4}_{i}\right)\exp\left(\theta_{i}+\theta_{n-i+1}\right)$
$=\frac{1}{2}\sum^{M}_{n=1}\sum^{n}_{i=1}a_{i}a_{n-i+1}\left(k^{2}_{n-i+1}-k^{2}_{i}\right)^{2}\exp\left(\theta_{i}+\theta_{n-i+1}\right).$
Under the hypotheses of the lemma, in general there are $\frac{1}{2}\left(M^{2}-M\right)$ linearly independent terms in the last expression, proving the required inclusion in $\mathcal{W}_{\frac{1}{2}M\left(M-1\right)}$. The $M$ terms correspond to the diagonal, and $M^{2}$ represents all the possible elements in an $M\times M$ matrix. ∎
@@ 3. $(iii)$
$\quad=-3k^{3}_{i}k_{j}+3k_{i}k^{3}_{j}=3\left(k_{i}k^{3}_{j}-k^{3}_{i}k_{j}\right)=3k_{i}k_{j}\left(k^{2}_{j}-k^{2}_{i}\right).$
@@
This proves that $Ai\left(\Theta\right)\in\mathcal{W}_{4}$.222 Assume $k_{1}<k_{2}<k_{3}<k_{4}$. Let us study the condition $Ai\left(\Theta\right)=0$. If $k_{1}=0$, then $k_{2}=-k_{3}=-k_{4}$, which is impossible. A similar argument holds if now $k_{2}=0$, or $k_{3}=0$, or $k_{4}=0$. Therefore, each $k_{i}$ must be nonzero. However, in this case $k_{1}=-k_{3}=-k_{4}$, also impossible. Thus, $Ai\left(\Theta\right)\neq 0$ for a nondegenerate 2-soliton. Replacing the values of the derivatives in $H\left(\Theta\right)$,
@@
Now, replacing the values of the derivatives in $W_{y}\left(\Theta\right)$,
@@ 5. Proof of Main Results
Now, Lemma 3.5 implies that $W_{y}\left(\Theta\right)-W_{x}\left(\Theta\right)=0$. Since $F=\log$, $W_{y}\left(\Theta\right)=W_{x}\left(\Theta\right)=0$. Since $W_{y}\left(\Theta\right)=W_{x}\left(\Theta\right)=Ai\left(\Theta\right)=0$, Lemma 3.11 ensures that $\Theta=A\left(t,x\right)\exp\left(cy\right)$, with $A\left(t,x\right)$ being of Airy type and with $c\in\mathbb{R}$.
@@
$A\left(t,x\right)=c_{A,1}\left(t\right)\left(t\right)\exp\left(kx\right)+c_{A,2}\left(t\right)\left(t\right)\exp\left(-kx\right),\quad k=\sqrt{c}\in\mathbb{C}-\{0\}.$ (5.2)
@@
$\Theta^{2}\partial^{2}_{y}\left(\ln\left(\Theta\right)\right)=A\left(t,x\right)\exp\left(k\left(t,x\right)y\right).$
This equation can be studied like a nonlinear second order ODE on $\Theta$ and the variable $y$. Considering the change of variable $f=\ln\left(\Theta\right)$ one gets
@@
From Corollary 4.8 we know that $u$ solution of (1.1) of the form (1.2) with smooth real-valued phase $\Theta>0$ corresponding to an $M$ resonant multi-soliton (1.20) and $F=\log$ imply $H\left(\Theta\right)=Ai\left(\Theta\right)=0$. Also, Lemma 4.7 in this particular case $\left(k_{j}^{2}\geq 0\right)$ states that $\Theta W_{y}\left(\Theta\right)=\Theta W_{x}\left(\Theta\right)\in\mathcal{W}_{\frac{1}{2}M\left(M-1\right)}$.
@@
Recall that each $b_{j}>0$. After arranging terms as members of an upper triangular matrix, we get
@@
Now we improve the coefficients using that $\Theta$ satisfies zero Heat and Airy, showing that $\Theta$ is a line-multi-soliton. Since Heat and Airy are linear equations, the proof is similar to the proof of Lemma 5.2, Step 2. More precisely,
@@
$\leavevmode\nobreak\ {}+\sum_{j=1}^{M}\left\{\left(a_{1,j,t}-a_{1,j}\sqrt{k^{3}_{j}}\right)-\frac{a_{1,j}k_{j,t}}{2\sqrt{k_{j}}}x+a_{1,j}k_{j,t}y\right\}\exp\left(-\sqrt{k_{j}}x+k_{j}y\right)$
@@
$a_{0,j}=a_{0,0,j}\exp\left(\sqrt{k^{3}_{j}}t\right),\quad a_{1,j}=a_{1,0,j}\exp\left(\sqrt{k^{3}_{j}}t\right),$
@@
$\Theta=\sum_{j=1}^{M}\left(a_{0,0,j}\exp\left(-k_{j}x\right)+\exp a_{1,0,j}\exp\left(k_{j}x\right)\right)\exp\left(k_{j}^{2}y+k^{3}_{j}t\right).$
@@
has solution $u=0.$ Assume $x>0$. Thanks to the exponential bound on $x$ for $\Theta$, this is obtained by simply taking Laplace transform, solving the corresponding obtained ODE, and using the uniqueness of the inverse Laplace transform in the exponentially growing class of solutions. The remaining case $x<0$ is proved similarly by changing $u(t,x)$ by $u(t,-x)$. ∎
@@ 6. Proof in the ZK case
$\leavevmode\nobreak\ {}+\sum_{j=2}^{d}\left(\Theta_{x_{j}x_{j}}F^{\prime}(\Theta)+\Theta_{x_{j}}^{2}F^{\prime\prime}(\Theta)\right)_{xx}.$
The term $\Theta_{x_{j}x_{j}}F^{\prime}(\Theta)+\Theta_{x_{j}}^{2}F^{\prime\prime}(\Theta)$ can be written as
$F^{\prime}(\Theta)W_{x_{j}}^{F}(\Theta)+\Theta_{x_{j}}^{2}\rho(\Theta),\quad\rho=F^{\prime\prime}(\Theta)+F^{\prime 2}(\Theta).$
Therefore, following the same ideas as in the KP-II case, we arrive to the equation
$0=$ $\leavevmode\nobreak\ {}\text{EDO}_{4}(\rho)-F^{\prime}(\Theta)(Ai(\Theta))_{x}+F^{\prime 2}(\Theta)(3W_{1}(\Theta)+\Theta_{x}Ai(\Theta))+\sum_{j=2}^{d}\left(F^{\prime}(\Theta)W_{x_{j}}^{F}(\Theta)\right)_{xx},$
where $\text{EDO}_{4}(\rho)$ is a fourth order linear ODE for $\rho$. Now we finish the proof. If $u$ is a KdV soliton and $F=\log$, it is easy to check that (1.23) is satisfied. Now, if (1.23) is satisfied, $\text{EDO}_{4}(\rho)=0$ and $F=\log$. Therefore, $W_{x_{j}}^{F}(\Theta)=W_{x_{j}}^{\log}(\Theta)=0$. From $W_{1}(\Theta)=0$ we get
$\Theta=a(x_{2},\ldots,x_{d},t)\exp\left(xb(x_{2},\ldots,x_{d},t)\right)+c(x_{2},\ldots,x_{d},t).$
From $W_{x_{j}}^{\log}(\Theta)=0$ we get that $a,b,c$ are only time dependent. Finally, from Airy we obtain that $u=Q_{k}$, for some constant $k>0$.

### v2

@@ 3. $\left(iii\right)$
The proof of $\left(i\right)$ is a consequence of the following identity. One has
@@ 3.1. Simple Phases
As in the proof of Lemma 3.7, we only prove the necessary condition. Thanks to this last result, $\Theta\left(t,x,y\right)=A\left(t,x\right)\exp\left(c\left(t,x\right)y\right)$. Assuming now $W_{x}\left(\Theta\right)=0$, i.e. $\Theta\Theta_{xxxx}-\Theta^{2}_{xx}=0$, and replacing $\Theta$ in this equation, one gets the next degree 3 polynomial in the variable $y$:
@@ 4.1. Airy and Heat structures
In the case $c_{2}\left(t\right)=0$ we first have $\Theta\left(t,x,y\right)=A\left(t,x\right)\exp\left(c_{1}\left(t\right)y\right)$. Since $H\left(\Theta\right)=0$ necessarily $A\left(t,x\right)c_{1}\left(t\right)-A_{xx}\left(t,x\right)=0$. Depending on the sign of $c_{1}\left(t\right)$, we have
$A\left(t,x\right)=a_{0}\left(t\right)e^{\sqrt{c_{1}\left(t\right)}x}+a_{1}\left(t\right)e^{-\sqrt{c_{1}\left(t\right)}x},\quad c_{1}\left(t\right)>0,$
$A\left(t,x\right)=a_{0}\left(t\right)+a_{1}\left(t\right)x,\quad c_{1}\left(t\right)=0,$
@@
$A\left(t,x\right)=a_{0}\left(t\right)\cos\left(\sqrt{-c_{1}\left(t\right)}x\right)+a_{1}\left(t\right)\sin\left(\sqrt{-c_{1}\left(t\right)}x\right),\quad-c_{1}\left(t\right)>0.$
Since $\Theta>0$ we discard the third case, and in the second case we get $c_{2}\left(t\right)=0$, leading to a solution included in the first case. Note that naturally the condition $A$ is of $W_{x}$ type is satisfied. Denoting $k_{1}:=\sqrt{c_{1}}$ we conclude. This proves (i).
@@
$\Theta=A\left(t,x\right)\exp\left(\left(c_{1}\left(t\right)+c_{2}\left(t\right)x\right)y\right).$
@@
Case $c_{2}\left(t\right)=0$. First, we will examine the case in which $c_{2}\left(t\right)=0$. In this case from Corollary 3.9 $\Theta=A\left(t,x\right)\exp\left(c_{1}\left(t\right)y\right)$, with $W_{x}\left(A\right)=0$. First we compute
@@
$\Theta\left(t,x,y\right)=a_{1}\left(t\right)\exp\left(k_{1}x+k_{1}^{2}y\right)+a_{2}\left(t\right)\exp\left(-k_{1}x+k_{1}^{2}y\right).$
@@
$\left(a_{1}^{\prime}-a_{1}k_{1}^{3}\right)\exp\left(k_{1}x+k_{1}^{2}y\right)+\left(a_{2}^{\prime}+a_{2}k_{1}^{3}\right)\exp\left(-k_{1}x+k_{1}^{2}y\right)=0.$
We obtain $a_{1}\left(t\right)=a_{1}e^{k_{1}^{3}t}$, $a_{2}\left(t\right)=a_{2}e^{-k_{1}^{3}t}$, where $a_{1}$ and $a_{2}$ are constants. Therefore,
@@
${\color[rgb]{0,0,0}\Theta\left(t,x,y\right)=\exp\left(y\partial_{x}^{2}\right)\Theta_{0}\left(t,x\right)=\frac{1}{\left(4\pi y\right)^{1/2}}\int_{\mathbb{R}}\exp\left(-\frac{\left(x-s\right)^{2}}{4y}\right)\Theta_{0}\left(t,s\right)\mathrm{d}s}$
@@ 4.2. Soliton structures
where $a_{1},a_{2}>0$, and $\theta_{j}:=k_{j}x+k_{j}^{2}y+k_{j}^{3}t$, $k_{1},k_{2}\in\mathbb{R}$. We shall say that $\Theta$ represents a line-soliton if $\Theta$ has the previous form. Notice that unless $F=\log$, one does not have that $u=2\partial_{x}^{2}F\left(\Theta\right)$ is the standard KP line-soliton (1.13).
@@
Assume that $\Theta\in\mathcal{W}_{M}$, that is for $0\leq k_{1}\left(t,x\right)<k_{2}\left(t,x\right)<\ldots<{\color[rgb]{0,0,0}k_{M}}\left(t,x\right)$,
@@
Note now that for each real numbers $A_{i},B_{j}$ with $i,j\in\{1,\cdots,M\}$ we have
$\sum^{M}_{n=1}A_{n}\sum^{M}_{n=1}B_{n}=\sum^{2M}_{n=2}C_{n}\hbox{, where }C_{n}=\sum_{\begin{subarray}{c}1\leq i,j\leq M\\ i+j=n\end{subarray}}A_{i}B_{j}.$
Thus
@@
$=\sum^{2M}_{n=2}\sum_{\begin{subarray}{c}1\leq i,j\leq M\\ i+j=n\end{subarray}}k^{4}_{j}a_{i}a_{j}\exp\left(\theta_{i}+\theta_{j}\right)$
$\quad-\sum^{2M}_{n=2}\sum_{\begin{subarray}{c}1\leq i,j\leq M\\ i+j=n\end{subarray}}k^{2}_{i}k^{2}_{j}a_{i}a_{j}\exp\left(\theta_{i}+\theta_{j}\right)$
$=\sum^{2M}_{n=2}\sum_{\begin{subarray}{c}1\leq i,j\leq M\\ i+j=n\end{subarray}}a_{i}a_{j}\left(k^{4}_{j}-k^{2}_{i}k^{2}_{j}\right)\exp\left(\theta_{i}+\theta_{j}\right).$
We observe that every diagonal contribution, where the two indices coincide, the factor $k_{i}^{4}-k_{i}^{2}k_{i}^{2}$ is identically zero. For the remaining off-diagonal terms, using $\theta_{i}+\theta_{j}=\theta_{j}+\theta_{i}$, we can pair the summands that come from $(i,j)$ and $(j,i)$. Adding those two contributions, associated to $\exp\left(\theta_{i}+\theta_{j}\right)$, replaces the original coefficient $k_{j}^{4}-k_{i}^{2}k_{j}^{2}$ by the symmetric combination $\left(k_{i}^{2}-k_{j}^{2}\right)^{2}$. After this last expresión reduces to a single sum obtaining
$\Theta W_{y}\left(\Theta\right)=\sum_{1\leq i<j\leq M}a_{i}a_{j}\left(k_{i}^{2}-k_{j}^{2}\right)^{2}e^{\theta_{i}+\theta_{j}}.$
Thus under the hypotheses of the lemma, are $\frac{1}{2}M(M-1)$ linearly independent terms in the last expression, proving the required inclusion in $\mathcal{W}_{\frac{1}{2}M\left(M-1\right)}$.
@@ 3. $(iii)$
$\quad=-3k^{3}_{i}k_{j}+3k_{i}k^{3}_{j}$
$\quad=3k_{i}k_{j}\left(k^{2}_{j}-k^{2}_{i}\right).$
@@
This proves that $Ai\left(\Theta\right)\in\mathcal{W}_{4}$.111 Assume $k_{1}<k_{2}<k_{3}<k_{4}$. Let us study the condition $Ai\left(\Theta\right)=0$. If $k_{1}=0$, then $k_{2}=-k_{3}=-k_{4}$, which is impossible. A similar argument holds if now $k_{2}=0$, or $k_{3}=0$, or $k_{4}=0$. Therefore, each $k_{i}$ must be nonzero. However, in this case $k_{1}=-k_{3}=-k_{4}$, also impossible. Thus, $Ai\left(\Theta\right)\neq 0$ for a nondegenerate 2-soliton. Replacing the values of the derivatives in $H\left(\Theta\right)$,
@@
Now, replacing the values of the derivatives in $W_{y}\left(\Theta\right)$, it follows
@@ 5. Proof of Main Results
Now, Lemma 3.5 implies that $W^{F}_{y}\left(\Theta\right)-W^{F}_{x}\left(\Theta\right)=W_{y}\left(\Theta\right)-W_{x}\left(\Theta\right)=0$. Since $F=\log$, $W_{y}\left(\Theta\right)=W_{x}\left(\Theta\right)=0$. Since $W_{y}\left(\Theta\right)=W_{x}\left(\Theta\right)=Ai\left(\Theta\right)=0$, Lemma 3.11 ensures that $\Theta=A\left(t,x\right)\exp\left(cy\right)$, with $A\left(t,x\right)$ being of Airy type and with $c\in\mathbb{R}$.
@@
$A\left(t,x\right)=c_{A,1}\left(t\right)\exp\left(kx\right)+c_{A,2}\left(t\right)\exp\left(-kx\right),\quad k=\sqrt{c}\in\mathbb{C}-\{0\}.$ (5.2)
@@
$\Theta^{2}\partial^{2}_{y}\left(\log\left(\Theta\right)\right)=A\left(t,x\right)\exp\left(k\left(t,x\right)y\right).$
This equation can be studied like a nonlinear second order ODE for $\Theta$ in the variable $y$. Considering the change of variable $f=\log\Theta$ one gets
@@
Let $u$ be a (smooth) solution of (1.1) of the form (1.2) with a smooth real-valued phase $\Theta>0$ satisfying for $k=0,1,2,3$, the hypotheses (1.21):
$\partial_{x}^{k}\Theta\left(t,0,0\right),\quad\partial_{x}^{k}\partial_{y}\Theta\left(t,0,0\right)\quad\hbox{uniquely prescribed.}$.
Assume that the smooth profile $F$ satisfies (1.17). Assume that $\Theta$ corresponds to an $M$ resonant multi-soliton (1.20) and $F=\log$. Since $\Theta>0$ as in (4.5) generates an $M$ resonant soliton, then from Corollary 4.8 $Ai\left(\Theta\right)=H\left(\Theta\right)=0$. Also, Lemma 4.7 in this particular case $\left(k_{j}^{2}\geq 0\right)$ states that $\Theta W_{y}\left(\Theta\right)=\Theta W_{x}\left(\Theta\right)\in\mathcal{W}_{\frac{1}{2}M\left(M-1\right)}$.
@@
Recall that each $b_{j}>0$. There are $\frac{1}{2}M\left(M-1\right)$ terms in the summation above. Relabeling and arranging terms above as members of a $M\times M$ lower triangular matrix, we get
@@
Now we improve the coefficients using that $\Theta$ satisfies zero Heat and Airy, showing that $\Theta$ is a multi-line-soliton. Since Heat and Airy are linear equations, the proof is similar to the proof of Lemma 5.2, Step 2. More precisely,
@@
$\leavevmode\nobreak\ {}+\sum_{j=1}^{M}\left\{\left(a_{1,j,t}+a_{1,j}\sqrt{k^{3}_{j}}\right)-\frac{a_{1,j}k_{j,t}}{2\sqrt{k_{j}}}x+a_{1,j}k_{j,t}y\right\}\exp\left(-\sqrt{k_{j}}x+k_{j}y\right)$
@@
$a_{0,j}=a_{0,0,j}\exp\left(\sqrt{k^{3}_{j}}t\right),\quad a_{1,j}=a_{1,0,j}\exp\left(-\sqrt{k^{3}_{j}}t\right),$
@@
$\Theta=\sum_{j=1}^{M}\left(a_{0,0,j}\exp\left(k_{j}x+k^{3}_{j}t\right)+a_{1,0,j}\exp\left(-k_{j}x-k^{3}_{j}t\right)\right)\exp\left(k_{j}^{2}y\right).$
@@
has solution $u=0.$ Assume $x>0$. Thanks to the exponential bound on $x$ for $\Theta$, this is obtained by simply taking the Laplace transform, solving the corresponding obtained ODE, and using the uniqueness of the inverse Laplace transform in the exponentially growing class of solutions. The remaining case $x<0$ is proved similarly by changing $u(t,x)$ to $u(t,-x)$. ∎
@@ 6. Proof in the ZK case
$\leavevmode\nobreak\ {}+\sum_{j=2}^{d}\left(\Theta_{x_{j}x_{j}}F^{\prime}\left(\Theta\right)+\Theta_{x_{j}}^{2}F^{\prime\prime}\left(\Theta\right)\right)_{xx}.$
The term $\Theta_{x_{j}x_{j}}F^{\prime}\left(\Theta\right)+\Theta_{x_{j}}^{2}F^{\prime\prime}\left(\Theta\right)$ can be written as
$F^{\prime}\left(\Theta\right)W_{x_{j}}^{F}\left(\Theta\right)+\Theta_{x_{j}}^{2}\rho\left(\Theta\right),\quad\rho=F^{\prime\prime}\left(\Theta\right)+F^{\prime 2}\left(\Theta\right).$
Therefore, following the same ideas as in the KP case, we arrive to the equation
$0=$ $\leavevmode\nobreak\ {}\text{EDO}_{4}\left(\rho\right)-F^{\prime}\left(\Theta\right)\left(Ai\left(\Theta\right)\right)_{x}$
$\leavevmode\nobreak\ {}+F^{\prime 2}\left(\Theta\right)\left(3W_{1}\left(\Theta\right)+\Theta_{x}Ai\left(\Theta\right)\right)+\sum_{j=2}^{d}\left(F^{\prime}\left(\Theta\right)W_{x_{j}}^{F}\left(\Theta\right)\right)_{xx},$
where $\text{EDO}_{4}\left(\rho\right)$ is a fourth order linear ODE for $\rho$. Now we finish the proof. If $u$ is a KdV soliton and $F=\log$, it is easy to check that (1.23) is satisfied. Now, if (1.23) is satisfied, $\text{EDO}_{4}\left(\rho\right)=0$ and $F=\log$. Therefore, $W_{x_{j}}^{F}\left(\Theta\right)=W_{x_{j}}^{\log}\left(\Theta\right)=0$. From $W_{1}\left(\Theta\right)=0$ we get
$\Theta=a\left(x_{2},\ldots,x_{d},t\right)\exp\left(xb\left(x_{2},\ldots,x_{d},t\right)\right)+c\left(x_{2},\ldots,x_{d},t\right).$
From $W_{x_{j}}^{\log}\left(\Theta\right)=0$ we get that $a,b,c$ are only time dependent. Finally, from Airy we obtain that $u=Q_{k}$, for some constant $k>0$.